GPS (Global Positioning System) satellites orbit at an altitude of 2.8×107 m . Part A. Find the orbital period. Express your answer using two significant figures. Part B Find the orbital speed of such a satellite. Express your answer using two significant figures.
Answer #1

mv^2/R = Gm*M/R^2

R = h + Ro

Ro = earth radius = 6400 km = 6.4 x 10^6 m

h = 2.8 x 10^7 m

R = 3.44 x 10^7 m

v^2 = GM/R

G = 6.67 x 10^-11

M = mass of earth = 6 x 10^24 kg

v^2 = GM/R

v = sqrt(GM/R)

v = 3.4 x 10^3 m/s

v = 2pi*r/T

T = 2pi*r/v

T = 6.34 x 10^4 s = 17.6 hrs = 18 hrs

part b )

v = 3.4 x 10^3 m/s